OpenStudy (ttop0816):

Please Help!! Q: You win a game if you roll a die and get a 2 or a 5. You play this game 60 times. What is the probability that you will win the game at least 15 times? If you can, show your work please!

OpenStudy (ttop0816):

@mathmale

OpenStudy (ttop0816):

@AAbomosalam1998 please help!

OpenStudy (kropot72):

Ths can be solved by using the normal approximation to the binomial distribution. The mean of the normal distribution is np, where n is the number of tests and p is the probability of success. In this case np = 60 * 1/3 = 20. The standard deviation of the normal distribution is found from: \[\large SD=\sqrt{np(1-p)}=3.65\]